OpenStudy (anonymous):

if the point a (pi/6,-2) lies on the graph of f(x) =sin(x+pi)-d then the value of d is?

OpenStudy (anonymous):

So you have y = sin ( x + pi) - d. Now plug in your point and solve for d. - 2 = sin (pi/6 + pi) - d -2 = sin 7pi/6 - d d = 3/2

OpenStudy (anonymous):

isint pi over six plus pi equal to 2 pi over 6

OpenStudy (anonymous):

nvm i got it thanks\